# Exploring End-to-end Differentiable Neural Charged Particle Tracking – A Loss Landscape Perspective

### v1

@@ Abstract
Measurement and analysis of high energetic particles for scientific, medical or industrial applications is a complex procedure, requiring the design of sophisticated detector and data processing systems. The development of adaptive and differentiable software pipelines using a combination of conventional and machine learning algorithms is therefore getting ever more important to optimize and operate the system efficiently while maintaining end-to-end (E2E) differentiability. We propose for the application of charged particle tracking an E2E differentiable decision-focused learning scheme using graph neural networks with combinatorial components solving a linear assignment problem for each detector layer. We demonstrate empirically that including differentiable variations of discrete assignment operations allows for efficient network optimization, working better or on par with approaches that lack E2E differentiability. In additional studies, we dive deeper into the optimization process and provide further insights from a loss landscape perspective. We demonstrate that while both methods converge into similar performing, globally well-connected regions, they suffer under substantial predictive instability across initialization and optimization methods, which can have unpredictable consequences on the performance of downstream tasks such as image reconstruction. We also point out a dependency between the interpolation factor of the gradient estimator and the prediction stability of the model, suggesting the choice of sufficiently small values. Given the strong global connectivity of learned solutions and the excellent training performance, we argue that E2E differentiability provides, besides the general availability of gradient information, an important tool for robust particle tracking to mitigate prediction instabilities by favoring solutions that perform well on downstream tasks.
@@ 1 Introduction
Charged particle tracking (Strandlie & Frühwirth, 2010) is a central element in analyzing readouts generated by ionizing energy losses in particle detectors, with the goal of reconstructing distinct sets of coherent particle tracks from a discrete particle cloud measured over subsequent detector layers. Current state-of-the-art approaches in particle tracking heavily utilizes geometric deep learning in a two-phase predict-then-optimize approach (Shlomi et al., 2021; Duarte & Vlimant, 2022; Thais et al., 2022). In an initial prediction step, relations between detector readouts described as a graph structure are modelled and quantified. Then the predicted relationships are used in a separate and disconnected step to parameterize a discrete optimization problem generating disconnected track candidates. Separating the initial scoring from the final assignment operations, however, bares the risk of learning suboptimal solutions, that only minimize the intermediate optimization quantity of the scoring task (prediction loss) instead of the final task loss (Mandi et al., 2023). Furthermore, it prevents the integration of particle tracking in any other combined gradient-based optimization of any kind. This could include for example the joint optimization of reconstruction loss and other auxiliary losses, generated for a downstream task, as well as the integration of particle tracking in optimization pipelines for experiment design and optimization (Dorigo et al., 2023; Aehle et al., 2023b). However, providing end-to-end differentiable solutions is not trivial as the final track assignment involves discrete, piece-wise constant assignment operations where the gradient is undefined. The decision-focused learning or predict-and-optimize paradigm (Kotary et al., 2021; Mandi et al., 2023) provides a general framework where prediction of intermediate scoring and optimization for structured outputs are directly integrated in the training procedure by embedding the combinatorial solver as a component of the network architecture, allowing to optimize the main objective in an E2E manner. This framework has already been proven highly efficient for various fields of applications such as natural language processing, computer vision or planning and scheduling tasks (Mandi et al., 2023). In this work, we propose and explore a predict-and-optimize framework for charged particle tracking, providing to our knowledge the first fully differentiable charged particle tracking pipeline for high-energy physics [^1] The source code and data utilized in this study are detailed in Appendix E, including all relevant source code, datasets, and results necessary to reproduce the findings presented in this paper.. Our main contributions summarize as follows:
@@
    We reveal, analyzing the local and global structure of the loss landscape, strong connectivity between the trained representations both for E2E training and optimization of a prediction loss.
@@
    Further, we find significant predictive instabilities between the different training paradigms as well as across random initializations, arguing the importance of E2E differentiability for robust particle tracking, providing necessary tools to constraint the solution space of tracking models to a subset optimizing an additional metric of the downstream tasks.
@@ 2 Related Work
Charged particle tracking is a well-studied problem, with the majority of advancements originating from developments in high-energy physics research at particle facilities (e.g., CERN). For the last decades, tracking algorithms have been dominated by conventional model based, pattern recognition and optimization algorithms such as Kalman filter (Frühwirth, 1987), Hough transform (Hough, 1959) or cellular automaton (Glazov et al., 1993). Here we specifically want to point out existing work by Pusztaszeri et al. (1996) related to our work, using combinatorial optimization on handcrafted features for particle tracking providing solutions that satisfy unique combination of vertex detector observations using a five-dimensional assignment model. With the ever-accelerating advancement and availability of deep learning algorithms, coupled with the increasing complexity and density of collision events, a significant demand has been placed on novel reconstruction algorithms, resulting in faster and more precise models. While initial studies mainly relied on basic regression- and classification models based on convolutional neural networks (CNN) or recurrent neural networks (RNN) (Farrell et al., 2017a; b), remarkable progress has been made recently by computationally efficient and well-performing network architectures based on geometric deep learning (GDL) (Shlomi et al., 2021; Thais et al., 2022) leveraging sparse representations of particle events in combination with graph neural networks (GNN). Here, the vast majority of current designs can be categorized into one of two main schemes. Edge classification tracking (Farrell et al., 2018; Ju et al., 2020; Duarte & Vlimant, 2020; Baranov et al., 2019; Heintz et al., 2020; DeZoort et al., 2021; Elabd et al., 2022; Murnane et al., 2023) predicts for each graph edge connecting two particle hits in subsequent layers a continuous output score which is then used to construct track candidates. In object condensation tracking (Kieseler, 2020; Qasim et al., 2022; Lieret & DeZoort, 2023; Lieret et al., 2023), GNN’s are trained with a multi-objective object condensation loss. In contrast to scoring edges in the graph, object condensation embeds particle hits in an N-dimensional cluster space, where hits belonging to the same track are attracted to a close proximity while hits from different tracks are repelled from each other. While both approaches provide great tracking performances, they require a separate optimization step (e.g., clustering) to generate final track candidates in a disconnected step, thus lacking an end-to-end differentiable architecture, preventing gradient flow throughout the whole reconstruction process. Recent progress has been made using reinforcement learning (Kortus et al., 2023) for charged particle tracking, allowing to differentiate through the discrete assignment operation using variants of the policy gradient or score function estimator approach maximizing an objective function $J(\theta)$, defined as the expected future reward obtained by the agent’s policy.
@@ 3 Theory and Background
Figure 1: Simplified detector geometry of the Bergen pCT Detector used as a basis for the data simulation using Monte-Carlo simulation. Figure from Kortus et al. (2023) (CC BY 4.0).
In this work, we focus on reconstructing particle tracks generated in the pixelated proton computed tomography (pCT) detector prototype, proposed by the Bergen pCT Collaboration (Alme et al., 2020; Aehle et al., 2023a). This detector is designed for generating high-resolution computer tomographic images using accelerated protons and is composed using a total of 4644 high-resolution ALPIDE CMOS Pixel sensors (Mager, 2016; Aglieri Rinella, 2017) developed for the upgrade of the Inner Tracking System (ITS) of the ALICE experiment at CERN. To capture the residual energy and path of the particle, the detector is arranged in a multi-layer structure, aligned in two tracking and 41 calorimeter layers. Each of the tracking layers, used for capturing the incoming particle path with as little interaction as possible, is separated from each subsequent layer by a 55.18 mm and 39.58 mm air gap respectively. The following calorimeter layers are separated from each other with a 3.5 mm aluminum slate. A detailed description of the structure and the material budget of the detector prototype is depicted in Figure 1. In this detector setup, multiple simultaneous particle trajectories, generated by a 230 MeV scanning pencil beam, are captured distal to a patient as discrete readouts of pixel clusters in the sensitive layers which have to be reconstructed into particle tracks under the influence of physical interactions with the detector material.
@@ 3.1 Particle Interactions in Matter
The path of protons at energies relevant for pCT is mainly influenced by interactions of the particles with both electrons and atomic nuclei changing the path of the accelerated particle (Gottschalk, 2018; Groom & Klein, 2000), baring additional difficulties in recovering the traversal path of the particle throughout a multi-layer particle detector. In the following, we briefly describe all three relevant effects in increasing order regarding their influence on the particle trajectories:
@@
    Ionizing energy loss: When interacting with negatively charged electrons in the atom’s outer shell, protons lose a fraction of their initial energy, due to attracting forces between the particles (Bloch, 1933). Due to the relatively small mass of the electron compared to the high momentum of the accelerated proton, the proton remains on its original path.
@@ 3.2 Loss Landscape Analysis
[…]
@@
Analyzing network similarities in both representations and outputs is an effective tool in comparing results of network configurations, providing an estimate of proximity in the loss landscape, especially for globally well-connected minima (Yang et al., 2021). Achieving high network similarity is especially desirable for our use cases, as differences in reconstructed tracks might influence results of the downstream tasks, such as image reconstruction for pCT or statistical analysis for HEP experiments. To gain an insight into the behavior of the trained networks, we thus quantify both the similarity of learned representations using CKA similarities (Kornblith et al., 2019) and prediction instability using the min-max normalized disagreement (Klabunde et al., 2023).
@@
    $d(f_{1},f2)=\mathbb{E}_{x,f_{1},f_{2}}1\left\{\arg\max f_{1}(\bm{x})\neq\arg\max f_{2}(\bm{x})\right\}.$ (5)
@@ 4.1 Hit Graph and Edge (Line) Graph Construction
In this work, we parameterize the detector data as an undirected line graph $\mathcal{G}_{L}=(\mathcal{V}_{L},\mathcal{E}_{L})$ generated from an initially generated hit graph $\mathcal{G}_{H}=(\mathcal{V}_{H},\mathcal{E}_{H})$ containing all the detector readouts of a single readout frame. In an initial hit graph structure, we represent all particle hit centroids in a readout frame as a set of graph vertices and describe possible track connections (segments) between two consecutive layers in the detector by edge connections. To capture a richer representation of the scattering behavior, providing a strong inductive bias for the message passing mechanism of graph neural networks, we transform this initial representation into a line graph ($\mathcal{G}_{L}$), where each edge in $\mathcal{G}_{H}$ is transformed into a node in $G_{L}$ (ref. Figure 2). Further, edges are generated between nodes $v_{L,i},v_{L,j}$ that share a common vertex, constructing descriptions of track deflections of candidates over three consecutive layers. This transformed representation allows to efficiently aggregate information of scattering behavior (two nodes connected by an edge in $\mathcal{G}_{H}$) providing important information on possible track segments in a single GNN message, as supposed to the more complex aggregation in $\mathcal{G}_{H}$ over a two-hop neighborhood. To provide machine-readable and differentiable features (w.r.t. simulation output) we parameterize both vertices ($\bm{v}_{i}$) and edges ($\bm{e}_{ij}$) as
$\begin{split}\bm{v}_{i}&=\left(\Delta E_{H,i}\|\Delta E_{H,j}\|\langle x,y,z\rangle_{H,i}\|\langle x,y,z\rangle_{H,j}\right)\\ \bm{e}_{ij}&=\begin{cases}\sin(\omega\cdot[1-\sqrt{s_{\cos}}])&\text{if }i=2k\\ \cos(\omega\cdot[1-\sqrt{s_{\cos}}])&\text{if }i=2k+1\\ \end{cases},\end{split}$ (8)
where $\Delta E_{H}$ is the deposited energy in the sensitive layers and $\langle x,y,z\rangle_{H}$ is the global position of the hit in the detector [^2] We use the continuous energy deposition and position outputs of the Monte-Carlo simulations here to be able to calculate gradients w.r.t. position and energy deposition. The parameters can be replaced with the cluster size and cluster centroid position accordingly. However, this representation requires additional work to provide differentiable solutions for pixel clustering. for the adjacent nodes $i$ and $j$. Further, each edge is parameterized by a positional encoding, similar to Vaswani et al. (2017). However, instead of using discrete token positions, we leverage cosine similarities $s_{\cos}$ between two hit segments as proposed by Kortus et al. (2023). To reduce the complexity of the graph and to minimize the combinatorial explosion of edges in $\mathcal{G}_{L}$, we remove edges with high angles in $\mathcal{G}_{H}$, measured orthogonal to the sensitive layer. We find suitable thresholds as a trade-off between reducing the graph size and minimizing the number of removed true edges (see Appendix A).
@@
Figure 2: Schematic representation of directed hit graph ($\mathcal{G}_{H}$) and undirected line graph ($\mathcal{G}_{L}:=L(\mathcal{G}_{H}$)) using detector readout data simulated for the Bergen pCT prototype detector over multiple detector layer.

### v2

@@ Abstract
Measurement and analysis of high energetic particles for scientific, medical or industrial applications is a complex procedure, requiring the design of sophisticated detector and data processing systems. The development of adaptive and differentiable software pipelines using a combination of conventional and machine learning algorithms is therefore getting ever more important to optimize and operate the system efficiently while maintaining end-to-end (E2E) differentiability. In this work, we lay the groundwork for E2E differentiable decision focused learning for the application of charged particle tracking using graph neural networks with combinatorial components, solving a linear assignment problem for each detector layer. We demonstrate empirically that including differentiable variations of discrete assignment operations allows for efficient network optimization, working better or on par with approaches that lack E2E differentiability. In additional studies, we dive deeper into the optimization process and provide further insights from a loss landscape perspective, providing a robust foundation for future work. We demonstrate that while both methods converge into similar performing, globally well-connected regions, they suffer under substantial predictive instability across initialization and optimization methods, which can have unpredictable consequences on the performance of downstream tasks such as image reconstruction. We also point out a dependency between the interpolation factor of the gradient estimator and the prediction stability of the model, suggesting the choice of sufficiently small values. Given the strong global connectivity of learned solutions and the excellent training performance, we argue that E2E differentiability provides, besides the general availability of gradient information, an important tool for robust particle tracking to mitigate prediction instabilities by favoring solutions that perform well on downstream tasks.
@@ 1 Introduction
Charged particle tracking (Strandlie & Frühwirth, 2010) is a central element in analyzing readouts generated by ionizing energy losses in particle detectors, with the goal of reconstructing distinct sets of coherent particle tracks from a discrete particle cloud measured over subsequent detector layers. Current state-of-the-art approaches in particle tracking heavily utilizes geometric deep learning in a two-phase predict-then-optimize approach (Shlomi et al., 2021; Duarte & Vlimant, 2022; Thais et al., 2022). In an initial prediction step, relations between detector readouts described as a graph structure are modeled and quantified. Then the predicted relationships are used in a separate and disconnected step to parameterize a discrete optimization problem, generating disconnected track candidates. Separating the initial scoring from the final assignment operations, however, bares the risk of learning suboptimal solutions, that only minimize the intermediate optimization quantity of the scoring task (prediction loss) instead of the final task loss (Mandi et al., 2023). This effect is potentially further magnified by integrating particle tracking in entire component-based data processing pipelines [^1] Processing pipelines for, e.g., Bergen pCT project and the ALICE, ATLAS and CMS experiments, depicting the complexity and variety of individual chained components, are described in detail in Aehle et al. (2023b); Buncic et al. (2015); Åkesson et al. (2005); Lange & the CMSCollaboration) (2011). used for reconstruction and analysis, introducing additional complexities due to non-monotonic error dependencies of components on the downstream reconstruction performance, as well as complex error propagation due to entanglement of individual processing steps (Sculley et al., 2015; Nushi et al., 2017) (further discussed in Section 3.2). Providing end-to-end differentiable pipelines, where all individual components allow the propagation of gradient information, could significantly improve the overall performance of reconstruction and analysis and allow the integration of gradient-based experiment design and optimization (Dorigo et al., 2023; Aehle et al., 2023b). Providing end-to-end differentiable solutions is, however, not trivial, as the structured nature of tracks requires discrete, piecewise constant assignment operations where the gradient is undefined. While recent work, aims to bootstrap the identification of tracks to downstream tasks (Stroud et al., 2024; Reuter et al., 2024), thereby achieving end-to-end optimization, the high combinatorial complexity of finding track candidates that satisfy the imposed unique assignment constraints directly, is a major challenge. The complexity of differentiating particle tracking and reconstruction pipelines in general is further examined and discussed in Aehle et al. (2023b). The decision-focused learning or predict-and-optimize paradigm (Kotary et al., 2021; Mandi et al., 2023) provides a general framework where prediction of intermediate scoring and optimization for structured outputs are directly integrated in the training procedure by embedding the combinatorial solver as a component of the network architecture, allowing to optimize the main objective in an E2E manner. This framework has already been proven highly efficient for various fields of applications such as natural language processing, computer vision, or planning and scheduling tasks (Mandi et al., 2023). In this work, we propose and explore a predict-and-optimize framework for charged particle tracking, providing, to our knowledge, the first fully differentiable charged particle tracking pipeline for high-energy physics [^2] The source code and data utilized in this study are detailed in Appendix F, including all relevant source code, datasets, and results necessary to reproduce the findings presented in this paper.. Further, we aim to provide first insights into the importance and usability of end-to-end particle tracking for fully differentiable reconstruction and analysis pipelines analyzing the loss landscape (e.g., sharpness/flatness of minima, connectivity of local minima, and similarity of learned models) of the tracking network. We choose this simplification over an analysis of an entire pipeline to reduce the overall complexity, introduced by other components, obscuring the influence of E2E tracking on the overall optimization performance. Further, E2E differentiability is still an active area of research for high-energy physics, with differentiability yet to be provided or improved for some components, making a meaningful and robust analysis infeasible. Our main contributions in this paper summarize as follows, serving as a foundation for further work on E2E charged particle tracking:
@@
    We find significant predictive instabilities between the different training paradigms as well as across random initializations, arguing the importance of E2E differentiability for robust particle tracking, providing necessary tools to constrain the solution space of tracking models to a subset optimizing an additional metric of the downstream tasks.
@@
    Further, we reveal, analyzing the local and global structure of the loss landscape, strong connectivity between the trained representations both for E2E training and optimization of a prediction loss. Together with the high predictive instability, we argue that the strong connectivity of diverse reconstruction policies enables efficient end-to-end optimization in reconstruction and analysis pipelines, favoring tracking solutions that provide beneficial performance on downstream tasks.
@@ 2 Related Work
Charged particle tracking is a well-studied problem, with the majority of advancements originating from developments in high-energy physics research at particle facilities (e.g., CERN). For the last decades, tracking algorithms have been dominated by conventional model-based pattern recognition and optimization algorithms such as Kalman filter (Frühwirth, 1987), Hough transform (Hough, 1959), or cellular automaton (Glazov et al., 1993). Here we specifically want to point out existing work by Pusztaszeri et al. (1996) related to our work, using combinatorial optimization on handcrafted features for particle tracking providing solutions that satisfy unique combination of vertex detector observations using a five-dimensional assignment model. With the ever-accelerating advancement and availability of deep learning algorithms, coupled with the increasing complexity and density of collision events, a significant demand has been imposed on novel reconstruction algorithms, resulting in faster and more precise models. While initial studies mainly relied on basic regression- and classification models based on convolutional neural networks (CNN) or recurrent neural networks (RNN) (Farrell et al., 2017a; b), remarkable progress has been made recently by computationally efficient and well-performing network architectures based on geometric deep learning (GDL) (Shlomi et al., 2021; Thais et al., 2022) leveraging sparse representations of particle events in combination with graph neural networks (GNN). Here, the vast majority of current designs can be categorized into one of two main schemes. Edge classification tracking (Farrell et al., 2018; Ju et al., 2020; Duarte & Vlimant, 2020; Baranov et al., 2019; Heintz et al., 2020; DeZoort et al., 2021; Elabd et al., 2022; Murnane et al., 2023) predicts for each graph edge connecting two particle hits in subsequent layers a continuous output score, which is then used to construct track candidates. In object condensation tracking (Kieseler, 2020; Qasim et al., 2022; Lieret & DeZoort, 2023; Lieret et al., 2023), GNN’s are trained with a multi-objective object condensation loss. In contrast to scoring edges in the graph, object condensation embeds particle hits in an N-dimensional cluster space, where hits belonging to the same track are attracted to a close proximity while hits from different tracks are repelled from each other. While both approaches provide great tracking performances, they require a separate optimization step (e.g., clustering) to generate final track candidates in a disconnected step, thus lacking an end-to-end differentiable architecture, preventing gradient flow throughout the whole reconstruction process. Recent progress has been made using reinforcement learning (Kortus et al., 2023) for charged particle tracking, allowing to differentiate through the discrete assignment operation using variants of the policy gradient or score function estimator approach, maximizing an objective function $J(\theta)$, defined as the expected future reward obtained by the agent’s policy. Additional work by e.g., Stroud et al. (2024); Reuter et al. (2024) proposed the fusion of tracking with track-fitting procedures (track parameter estimation), avoiding the optimization over discrete assignment operations.
@@ 3 Theory and Background
Figure 1: Schematic representation of the combinatorial complexity of reconstructing particle tracks over multiple subsequent layers. Marked in bold lines is the correct particle track. Marked in dashed lines are only possible next track segments originating from the previous correct segment. During reconstruction, all possible combinations of segments over all hits are considered.
In this work, we focus on reconstructing particle tracks generated in the pixelated proton computed tomography (pCT) detector prototype, proposed by the Bergen pCT Collaboration (Alme et al., 2020; Aehle et al., 2023a). This detector is designed for generating high-resolution computer tomographic images using accelerated protons and is composed using a total of 4644 high-resolution ALPIDE CMOS Pixel sensors (Mager, 2016; Aglieri Rinella, 2017) developed for the upgrade of the Inner Tracking System (ITS) of the ALICE experiment at CERN. To capture the residual energy and path of the particle, the detector is arranged in a multi-layer structure, aligned in two tracking and 41 calorimeter layers. Each of the tracking layers, used for capturing the incoming particle path with as little interaction as possible, is separated from each subsequent layer by a 55.18 mm and 39.58 mm air gap respectively. The following calorimeter layers are separated from each other with a 3.5 mm aluminum slate. A detailed description of the detector geometry can be found in Alme et al. (2020). In this detector setup, multiple simultaneous particle trajectories, generated by a 230 MeV scanning pencil beam, are captured distal to a patient as discrete readouts of pixel clusters in the sensitive layers which have to be reconstructed into particle tracks under the influence of physical interactions with the detector material, randomly deflecting the particle from its straight path (ref. Appendix C). With the high particle density in the detector, the resulting combinatorial explosion of track candidates, where each hit combination in subsequent layers need to be considered as possible partial solutions, requires efficient and reliable algorithms that are able to simultaneously reconstruct disjoint particle tracks by uniquely connecting particle hits over multiple layers in the detector (ref. Figure 1).
@@ 3.1 Particle Interactions in Matter
The path of protons at energies relevant for pCT is mainly influenced by interactions of the particles with both electrons and atomic nuclei changing the path of the accelerated particle (Gottschalk, 2018; Groom & Klein, 2000), posing additional difficulties in recovering the traversal path of the particle throughout a multi-layer particle detector. In the following, we briefly describe all three relevant effects in increasing order regarding their influence on the particle trajectories:
@@
    Ionizing energy loss: When interacting with negatively charged electrons in the atom’s outer shell, protons lose a fraction of their initial energy, due to attracting forces between the particles (Bloch, 1933), causing the particle to stop after a traversal range, proportional to the particle’s energy. Due to the relatively small mass of the electron compared to the high momentum of the accelerated proton, the proton remains on its original path.
@@
3.2 Component-based Processing Pipelines
Modern data processing pipelines are increasingly powered by data-driven or machine learning algorithms, requiring the efficient interplay of multiple individually optimized tasks that are strongly coupled and entangled with each other (Sculley et al., 2015; Nushi et al., 2017). The coupling of data-driven algorithms introduces additional challenges that need to be considered in order to ensure operability and maintainability while providing reproducible and optimal performance Sculley et al. (2015). While the influences of component-based processing pipelines utilizing machine learning are manyfold, we focus in the following on the following aspects, that influence the initial design and operation, especially benefiting from end-to-end optimizable/differentiable solutions:
  • Stochastic nature of optimization: In contrast to conventional model- or rule-based algorithms, data-driven algorithms are sensitive to random mechanisms, such as parameter initialization, stochastic optimization, or parallel operations on GPUs, reducing the overall stability of the system.
  • Error propagation: While individual model performances are often not only affected by prediction instabilities (Lopez-Martinez et al., 2022), the different sources of errors propagate throughout the entire pipeline, resulting in unpredictable error patterns in the final output of the pipeline due to the existent complex entanglements of the downstream processing steps.
  • Non-monotonic error: Finally, due to the entanglement of the individual components, model improvements of individual components are not necessarily bound to an improvement of the overall system performance Nushi et al. (2017). While this seems unintuitive at first, this behavior can be caused by downstream tasks being optimized on the erroneous output signal of the upstream part of the pipeline or misalignment of the individual components’ optimization metric and overall system performance.
To tackle the challenges associated with component-based pipelines, different solutions have been developed for mitigating the effect. This includes, for example, the development of modular processing pipelines, aiming to decouple the individual pipeline components to reduce the complex effects of entanglements Modi et al. (2023). Increasing effort has also been put into the development of E2E-differentiable or in general optimizable ML pipelines (Yu et al., 2022; Hilprecht et al., 2023), allowing to optimize the entirety of components while optimizing the final output of interest. A similar trend, utilizing end-to-end differentiable solutions, can also be found in high-energy physics (Dorigo et al., 2023; Aehle et al., 2023b).
@@ 3.3 Loss Landscape Analysis
$\mathcal{L}(\theta)=\int_{0}^{1}\mathcal{L}(\psi_{\theta}(t))dt=\mathbb{E}_{t\sim U(0,1)}\left[\mathcal{L}\left(\psi_{\theta}(t)\right)\right].$ (2)
@@
Analyzing network similarities in both representations and outputs is an effective tool in comparing results of network configurations, providing an estimate of proximity in the loss landscape, especially for globally well-connected minima (Yang et al., 2021).
Achieving high network similarity is especially desirable for our use cases, as differences in reconstructed tracks might influence results of the downstream tasks, such as image reconstruction for pCT or statistical analysis for HEP experiments.
To gain an insight into the behavior of the trained networks, we thus quantify both the similarity of learned representations using CKA similarities (Kornblith et al., 2019) and prediction instability using the min-max normalized disagreement (Klabunde et al., 2023).
@@
    $d(f_{1},f_{2})=\mathbb{E}_{x,f_{1},f_{2}}1\left\{\arg\max f_{1}(\bm{x})\neq\arg\max f_{2}(\bm{x})\right\}.$ (5)
@@ 4.1 Hit Graph and Edge (Line) Graph Construction
Figure 2: Schematic representation of directed hit graph ($\mathcal{G}_{H}$) and undirected line graph ($\mathcal{G}_{L}:=L(\mathcal{G}_{H}$)) using detector readout data simulated for the Bergen pCT prototype detector over multiple detector layers.
In this work, we parameterize the detector data as an undirected line graph $\mathcal{G}_{L}=(\mathcal{V}_{L},\mathcal{E}_{L})$ generated from an initially generated hit graph $\mathcal{G}_{H}=(\mathcal{V}_{H},\mathcal{E}_{H})$ containing all the detector readouts of a single readout frame. In an initial hit graph structure, we represent all particle hit centroids in a readout frame as a set of graph vertices $v_{H,i}$ and describe possible partial track connections (segments) between two consecutive layers in the detector by edge connections $e_{H,ij}$. This representation is widely used for particle tracking, as it allows describing and parameterizing the particle hits and possible partial solutions in a single sparse structure. Yet, to capture a richer representation of the scattering behavior over combinations of two consecutive segments, providing a strong inductive bias for modelling the most likely path of the particle in the detector, under the influence of elastic nuclear interactions, we transform this initial representation into a line graph ($\mathcal{G}_{L}$), where each edge in $\mathcal{G}_{H}$ is transformed into a node in $G_{L}$ (ref. Figure 2). Further, edges are generated between nodes $v_{L,i},v_{L,j}$ that share a common vertex in $\mathcal{G}_{H}$, constructing descriptions of track deflections of candidates over three consecutive layers. This transformed representation allows to efficiently aggregate and compare information of scattering behavior, providing important information on possible track segments in one edge feature, that can be aggregated by a single GNN message, as supposed to the more complex aggregation in $\mathcal{G}_{H}$ over a two-hop neighborhood (ref. Figure 2). To provide machine-readable and differentiable features (w.r.t. simulation output) we parameterize both vertex ($\bm{v}_{L,i}$) and edge ($\bm{e}_{L,ij}$) features [^3] In the following, unless otherwise specified, we use a simplified notation for $\bm{v}_{L,i}$ and $\bm{e}_{L,ij}$, where $\bm{v}_{i}$ and $\bm{e}_{ij}$ defaults to to edge- and node-features of the line graph representation $\mathcal{G}_{L}$. The additional identifyer $L$ is omitted in favor of readability. of the line-graph representation as
$\begin{split}\bm{v}_{L,i}&=\left(\Delta E_{H,i}\|\Delta E_{H,j}\|\langle x,y,z\rangle_{H,i}\|\langle x,y,z\rangle_{H,j}\right)\\ \bm{e}_{L,ij}&=\begin{cases}\sin(\omega\cdot[1-\sqrt{s_{\cos}}])&\text{if }i=2k\\ \cos(\omega\cdot[1-\sqrt{s_{\cos}}])&\text{if }i=2k+1\\ \end{cases},\end{split}$ (8)
where $\Delta E_{H}$ is the deposited energy in the sensitive layers, and $\langle x,y,z\rangle_{H}$ is the global position of the particle hit in the detector [^4] We use the continuous energy deposition and position outputs of the Monte-Carlo simulations here to be able to calculate gradients w.r.t. position and energy deposition. The parameters can be replaced with the cluster size and cluster centroid position accordingly. However, this representation requires additional work to provide differentiable solutions for pixel clustering. for the adjacent nodes $i$ and $j$. Further, each edge is parameterized by a positional encoding, similar to Vaswani et al. (2017). However, instead of using discrete token positions, we leverage cosine similarities $s_{\cos}$ between the directional vectors of two track segments in $\mathcal{G}_{H}$ as proposed by Kortus et al. (2023), injecting additional information on the relative likelihood of observing a specific transition dynamic of track candidates over segment combinations over three consecutive detector layers under the underlying physical interaction mechanisms. The importance and usefulness of the line-graph transformation, especially its beneficial parameterization of segment pairs, functioning as a strong inductive bias by condensing information on particle scattering into a single feature, is analyzed and confirmed in detail in Appendix D. Here, we highlight the performance gain compared to a conventional hit-graph baseline (DeZoort et al., 2021), parameterized according to the description in (Kortus et al., 2023). Additional results demonstrating the importance of segment information, provided by the positional encoding mechanism, are presented in (Kortus et al., 2023).
To reduce the complexity of the graph and to minimize the combinatorial explosion of edges in $\mathcal{G}_{L}$, we remove implausible edges with high angles in $\mathcal{G}_{H}$, measured orthogonal to the sensitive layer. We find suitable independent thresholds for both calorimeter and tracking layer edges as a trade-off between reducing the graph size and minimizing the number of removed true edges required for constructing the particle tracks (see Appendix A). Both thresholds are determined on the train-set and are kept constant for evaluation. Despite the elimination of graph edges, using a line-graph representation still introduces computational overhead for reconstruction. Additional results quantifying the impact of the line-graph representation are presented in Appendix D.